OpenStudy (anonymous):

I need help with answering the following problem: State the domain and range for the following. Find when f(x)=0. f(x)= (x+2)/(5x-10)

OpenStudy (itsonlycdeee):

Since f(x)=(x+2)/(5x+10) is the same thing as y=(x+2)/(5x+10) Domain = x and range = y. If you graph it, domain (negative infinity, negative 2) union (-2, positive infinity) and range is all real numbers.

OpenStudy (itsonlycdeee):

The domain is also except when x = -2 because that would make the denominator undefined.

OpenStudy (anonymous):

the domain is: 5x+10=0 x=-2 is impossible range i the domain of the inverse \[5xy+10y=x+2\rightarrow f^-1(x) \frac{ 2-10x }{5x-1 }\] so x=1/5 is impossible

OpenStudy (anonymous):

I.e the domain R-{-2} range R-{1/5}

OpenStudy (anonymous):

i got the domain part maybe you can explain the range more in depth? and why isn't the range just all real numbers?

OpenStudy (anonymous):

you have f(x) represented by the point in the plane (x,f(x)) the inverse when we solve for f(x) , I mean we try to find the point (f(x),x) ,so if we tried to find the entire domain of the inverse we get the range and that's what just happened

OpenStudy (anonymous):

anyways try to plug f(x)=1/5 and see \[\frac{ x+2 }{ 5x-10 }=1/5 \rightarrow x-2=x+2\] and that's impossible dude !!

OpenStudy (anonymous):

should i have mentioned that I'm a senior at a public school with 1 day of being in AP calc, explain it to me like i'm an idiot, because i'm still a calc virgin

OpenStudy (anonymous):

wait nvm i got it

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

How about f(x)=4/1-x ?

OpenStudy (anonymous):

lol even I didn't understand what you said about yourself , but I'm glad you got it after all :)